OpenStudy (anonymous):

Solve system by subsitution. 3x+y=22 -4x+8y=-20

OpenStudy (anonymous):

y= 22-3x ...-4x + 8(22-3x) = -20 solve for x then solve for y

OpenStudy (anonymous):

I really don't have a clue at all

OpenStudy (anonymous):

I haven't did this kind of math in over 30 years that why it strange and confusing

OpenStudy (anonymous):

3x+y=22 -4x+8y=-20 -8(3x+y=22) -24x-8y=-176 -4x+8y=-20 -28x=-198 -28 -28